OpenStudy (anonymous):

put this equation in y=a(x-h)^2+k or x=a(y-k)^2+h form:3(y-3)=(x+6)^2

OpenStudy (anonymous):

i kinda meant put 3(y-3)=(x+6)^2 in y=a(x-h)^2+k or x=a(y-k)^2+h form

OpenStudy (mathmate):

3(y-3)=(x+6)^2 y-3=(1/3)(x+6)^2 y=(1/3)(x+6)^2+3 a=(1/3) h=-6 k=3